OpenStudy (anonymous):

The tip of an 11-inch wiper blade wipes a path that is 31 inches long. What is the angle of rotation of the blade in radians to the nearest tenth? 5.6 radians 1.4 radians 0.4 radians 2.8 radians

OpenStudy (aaronq):

\(\theta=\dfrac{s}{r}=\dfrac{31}{11}\)

OpenStudy (anonymous):

?

OpenStudy (aaronq):

that is the answer?

OpenStudy (anonymous):

but its not a choice huh?

OpenStudy (aaronq):

i guess i have to spell it out, 31/11=2.8

OpenStudy (anonymous):

thank you sorry I just don't understand it

OpenStudy (anonymous):

bye then

OpenStudy (anonymous):

2.8 :)

OpenStudy (aaronq):